40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 29 21:50:21 2012, Jenne, Update, Locking, Yarm error spectra LSC_POY_11_I_ERR_29May2012.pdf
    Reply  Wed May 30 13:24:08 2012, Jenne, Update, Locking, Yarm error spectra 
Message ID: 6711     Entry time: Tue May 29 21:50:21 2012     Reply to this: 6714
Author: Jenne 
Type: Update 
Category: Locking 
Subject: Yarm error spectra 

The ~16Hz bounce mode of some optic is showing up in the Yarm error signal. 

MC is kind of 'windy' looking, so maybe it's from that? (Yuta's guess).

We need to make sure that the SUS damping and oplev paths both have notches at the correct bounce mode, not the old, old MOS frequency.  If that doesn't work, may need to put a resgain in Yarm path.

Attachment 1: LSC_POY_11_I_ERR_29May2012.pdf  19 kB  | Hide | Hide all
LSC_POY_11_I_ERR_29May2012.pdf
ELOG V3.1.3-